# Algebraic K-theory of finite algebras over higher local fields

### v2

@@ 1.1. Motivation and history
Algebraic K-theory is an important invariant for rings and even for ring spectra. It captures some arithmetic information of the input ring, and it reflects chromatic information of the input ring spectrum. Furthermore, it is not merely an invariant for ring spectra, but also a localizing invariant in the sense of [BGT13]. It also can output the sphere spectrum, which serves as the initial object in stable homotopy theory.
To be more precise about the information that algebra K-theory captures, we let $\mathcal{O}$ be a $\mathbb{Z}[1/\ell]$-algebra, where $\ell$ is a prime number. Quillen conjectured that there is a spectral sequence
@@
that converges to the target in large degrees. If moreover, $\mathcal{O}$ contains an $\ell^{n}$th root of unity, then Thomason proved that there is a spectral sequence:
@@
Here, the element $\beta\in\operatorname{K}_{2}(\mathcal{O},\mathbb{Z}/l^{n})$ corresponds to an $\ell^{n}$th root of unity. In this case, Bott inverted K-theory coincides with topological K-theory (defined in Thomason’s work); furthermore, in sufficiently large degrees, all three (Bott inverted, topological, and étale) become equivalent. It has been showed that the Bott inverted K-theory is equivalent to $L_{1}^{f}\operatorname{K}(\mathcal{O})\otimes\mathbb{S}/\ell\simeq L_{K(1)}\operatorname{K}(\mathcal{O})\otimes\mathbb{S}/\ell\simeq\operatorname{K}(\mathcal{O})\otimes\mathbb{S}/\ell[v_{1}^{-1}]$ for $\ell>2$. Moreover, when $\mathcal{O}$ does not contain $\ell^{n}$th roots of unity, there is a spectral sequence with the same $E_{2}$-page that converges to the étale K-theory.
@@
The original Lichtenbaum–Quillen conjecture is related to special values of the zeta function, a central topic in number theory. In contrast, Waldhausen’s reformulation of the conjecture adopts a more topological perspective, fitting naturally within chromatic homotopy theory. This contrast is striking. The K-theory of certain nice rings exhibit $v_{1}$-periodicity, because the localization is almost an equivalence. A natural question is whether there exists a higher level chromatic analogue. This question is addressed by Rognes’ red-shift conjecture. There are actually many forms of this conjecture, the key point of which it that, for a ring spectrum exhibiting $v_{n}$-periodicity, the K-theory of it should exhibit $v_{n+1}$-periodicity.
In studying these phenomena, we begin by fixing a prime number $p$. A spectrum $X$ is said to be fp spectrum, if it is $p$-complete, bounded below, and there is a $p$-local finite spectrum $F$ such that $F\otimes X$ is bounded, and the homotopy groups are finite groups. We say $X$ has fp-type at most $n$, if there is a type $n+1$ finite complex $F$, such that $F\otimes X$ is bounded, and the homotopy groups are finite. In his 2000 Oberwolfach lecture [Rog00], Rognes introduced the term ”redshift”, and the redshift problem is stated as follows.
@@
In chromatic homotopy theory, $v_{n}$ is detected by Morava K-theory, $K(n)$. One could use Morava K-theory to define the height for spectra, although, in general, this is not a very nice concept. We say an $\mathbb{E}_{\infty}$-ring $R$ is of height $n$, if $L_{K(n)}R\not\simeq 0$, and $L_{K(n+1)}R\simeq 0$. Hahn [Hah22] showed that for an $\mathbb{E}_{\infty}$-ring, if $L_{K(i)}R\simeq 0$, then $L_{K(j)}R\simeq 0$, for all $j\geq i$. Thus, in this case, the height is well defined. So, one might define the height to be the number $n$, such that $L_{K(i)}R\not\simeq 0$, for $0\leq i\leq n$, and $L_{K(n+1)}R\simeq 0$.
@@
The conjecture is true for $R$ any $\mathbb{E}_{\infty}$-ring, by [Yua21], [BSY22], [CMNN24]. For the relation of the first conjecture to $T(n)$-localized theory, see Theorem 4.5. The second conjecture is more related to $K(n)$-localized theory. As known, for $n\geq 2$, [BHLS23] shows the telescope conjecture fails, using many ingredients in algebraic K-theory and things related to these two conjectures.
@@
Let $R$ be an $\mathbb{E}_{1}$-ring spectrum of pure fp type $n$. Does $\operatorname{TR}(R,p)$ have pure fp type $n+1$.
The method to prove Conjecture 1.4, is first to construct a nice algebra map $R\to S$, yielding an algebra map $L_{K(n+1)}\operatorname{K}(R)\to L_{K(n+1)}\operatorname{K}(S)$. If one can show that $L_{K(n+1)}\operatorname{K}(S)\not\simeq 0$, then one could conclude that the source is not 00.
The K-theory of truncated polynomial algebras over perfect fields were studied in [HM97a]. Note that $\mathbb{F}_{p}[x]/x^{e}$ is just $\operatorname{BP}\langle-1\rangle[x]/x^{e}$. Thus it is natural to consider the K-theory of $\operatorname{BP}\langle n\rangle[x]/x^{e}$, which are chromatic higher level cases of $\mathbb{F}_{p}[x]/x^{e}$. The goal of this paper is to study Problem 1.2, Conjecture 1.3, 1.4, Question 1.5, 1.6, for $R=\operatorname{BP}\langle n\rangle[x]/x^{e}$, and other rings related to that. In particular, in the proof of Proposition 4.9 and [HM97a, Theorem A], shows that $\operatorname{K}(\mathbb{F}_{p}[x]/x^{e})$ is not an fp spectrum of type 00. And Proposition 6.4, shows that $\operatorname{K}((\ell[x]/x^{e})^{\wedge}_{p})^{\wedge}_{p}$ is not an fp spectrum of type $2$.
@@ 1.2. Main results
We prove the following.
@@
See Convention 2.12, 2.19. Any of them is a finite $\operatorname{BP}\langle n\rangle$-module for some $n$, thus we can view them as finite algebras over $\operatorname{BP}\langle n\rangle$ (the ring of integers of a height $n$ local field). Moreover $(\operatorname{BP}\langle n\rangle[x]/x^{e})^{\wedge}_{p},(\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e})^{\wedge}_{p}$ and $(\Lambda_{\operatorname{BP}\langle n\rangle}(x))^{\wedge}_{p},(\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\Lambda_{\operatorname{MU}}(x))^{\wedge}_{p}$ are fp spectra of type $n$, and the spectrum $(\operatorname{BP}\langle n\rangle/v_{n}^{e})^{\wedge}_{p}$ is an fp spectrum of type $n-1$.
@@
induces isomorphism on large degrees. In particular, if $F$ is of type $n+2$, the above also holds.
@@
$\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p\to L_{2}^{f}(\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p),$
$\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p\to L_{2}^{f}(\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])\otimes\mathbb{S}/p).$
One could say, even in the ring spectra case, algebraic K-theory still has the $\mathbb{A}^{1}$-invariance property.
By decomposing $\operatorname{THH}(\mathbb{S}[x]/x^{e})$ into a sum of spectra, and analyzing the Frobenius map, we found the following theorem. Actually, the base case is Proposition 4.9.
@@
Let $F$ be any type $n+2$ finite complex. Then the fiber of the Frobenius
@@
is not bounded above, where $e>1$. Moreover, $W\not\simeq 0$, where $W$ is the fiber of the Frobenius map of $F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)$. Consequently, $F\otimes\operatorname{TR}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is not bounded, where $e>1$. In other-words, the Lichtenbaum–Quillen property defined in Definition 1.7 fails for $\operatorname{BP}\langle n\rangle[x]/x^{e}$, when $e>1$.
@@
$\operatorname{THH}(\operatorname{MU}[x]/x^{e})\to\operatorname{THH}(\operatorname{MU}[x]/x^{e}),$
$\operatorname{THH}(\operatorname{BP}[x]/x^{e})\to\operatorname{THH}(\operatorname{BP}[x]/x^{e}),$
@@ 1.3. Future work
It seems that one could use the method presented in this paper to investigate some other examples of the Segal conjecture, by using the cyclic decomposition, i.e. decomposition of $\operatorname{THH}(\mathbb{S}[M])$ or $\operatorname{THH}(\mathbb{S}[\Omega Y])$, to carry out more computations, where $M$ is a commutative monoid, $Y$ is a space. For instance, in an forthcoming work, we try to compute $\operatorname{TC}(\operatorname{BP}\langle n\rangle[x,y]/(xy))$ for lower height cases. Similarly, on could replace $\operatorname{BP}\langle n\rangle$ by $K(n)$ or $k(n)=\tau_{\geq 0}K(n)$.
@@
Question 1.15 ([HLS24, Question 1.23]).
Does mod $(p,v_{1},\ldots,v_{n+1})$ syntomic cohomology for height $n$ ring spectra, factor though its mod $(p^{i_{0}},v_{i}^{i_{1}},\ldots,v_{n}^{i_{n}})$ reduction for some sequence $i_{0},\ldots,i_{n}$
@@ 1.4. Structure of the paper
In Section 2, we use several $\mathbb{E}_{\infty}$-ring maps to construct several $\mathbb{E}_{3}$-algebras over some base ring spectra other than $\mathbb{S}$, and we also use the tower of $\mathbb{E}_{\infty}$-algebras in [HW22, §2] to get an $\mathbb{E}_{3}$-algebra $\operatorname{BP}\langle n\rangle/v_{n}^{e}$. All these algebras are finite algebras over $\operatorname{BP}\langle n\rangle$, see Proposition 2.28. In Section 3, we use several $\mathbb{E}_{3}$-algebra maps to prove Conjecture 1.4, and to prove Theorem 1.10. In Section 4, we use the descent functor $\operatorname{desc}_{\mathbb{F}_{p}}$ to prove the Segal conjecture, and we use the decompositions of $\operatorname{THH}(\mathbb{S}[x]/x^{e}),\operatorname{THH}(\mathbb{S}[\Omega S^{3}])$ to prove the failure of the Segal conjecture for some ring spectra. The primary goal is to prove Theorem 1.11, 1.12, 1.13. In Section 5, we use the evenly free map $\operatorname{THH}(\operatorname{MU}[x])\to\operatorname{MU}[x]$ to compute the $E_{2}$-page of the descent spectral sequence which computes $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e};\mathbb{F}_{p}[x]/x^{e})$. We show that the spectral sequence degenerates at the $E_{2}$-page in some cases. See Proposition 5.4. In Section 6, we use the decomposition of $\operatorname{THH}(\mathbb{S}[x]/x^{e})$ try to carry out some computations for $V(2)_{*}\operatorname{TC}(\ell[x]/x^{e})$. We prove Proposition 6.4.

### v3

@@ 1.1. Motivation and history
Algebraic K-theory is an important invariant for rings and more generally ring spectra. It captures some arithmetic information of the input ring, and it reflects chromatic information of the input ring spectrum. Furthermore, it is not merely an invariant for ring spectra, but also an universal localizing invariant in the sense of [BGT13]. It also can output the sphere spectrum, which serves as the initial object in stable homotopy theory.
To be more precise about the information that algebra K-theory captures, we let $\mathcal{O}$ be a $\mathbb{Z}[1/\ell]$-algebra, where $\ell$ is a prime number. Quillen [QUI75] conjectured that there is a spectral sequence
@@
that converges to the target in large degrees. Lichtenbaum [LIC73] made a conjecture relate algebraic K-theory groups to special values of zeta functions. Wiles proved that the quotients of étale cohomology groups of $\mathcal{O}$ recover special values of zeta functions for certain nice rings [WIL90]. So, the conjecture made by Lichtenbaum coincides with Quillen’s conjecture for certain nice rings. The author learned this from [ROG99].
If moreover, $\mathcal{O}$ contains an $\ell^{n}$-th root of unity, then Thomason [THO85] proved that there is a spectral sequence:
@@
Here, the element $\beta\in\operatorname{K}_{2}(\mathcal{O},\mathbb{Z}/l^{n})$ corresponds to an $\ell^{n}$-th root of unity. In this case, Bott inverted K-theory coincides with topological K-theory (defined in Thomason’s work); furthermore, in sufficiently large degrees, all three (Bott inverted, topological, and étale) become equivalent. It has been showed that the Bott inverted K-theory is equivalent to $L_{1}^{f}\operatorname{K}(\mathcal{O})\otimes\mathbb{S}/\ell\simeq\operatorname{K}(\mathcal{O})\otimes\mathbb{S}/\ell[v_{1}^{-1}]\simeq L_{K(1)}\operatorname{K}(\mathcal{O})\otimes\mathbb{S}/\ell$ for $\ell>2$, see [THO85, Appendix A]. Moreover, when $\mathcal{O}$ does not contain $\ell^{n}$-th roots of unity, there is a spectral sequence with the same $E_{2}$-page that converges to the étale K-theory.
@@
The original Lichtenbaum–Quillen conjecture is related to special values of the zeta function, a central topic in number theory. In contrast, Waldhausen’s reformulation of the conjecture adopts a more topological perspective, fitting naturally within chromatic homotopy theory. The K-theory of certain nice rings exhibit $v_{1}$-periodicity, because the localization is almost an equivalence. A natural question is whether there exists a higher level chromatic analogue. This question is addressed by Rognes’ red-shift conjecture. There are actually many forms of this conjecture, the key point of which it that, for a ring spectrum exhibiting $v_{n}$-periodicity, the K-theory of it should exhibit $v_{n+1}$-periodicity.
In studying these phenomena, we begin by fixing a prime number $p$. A spectrum $X$ is said to be fp spectrum, if it is $p$-complete, bounded below, and there is a $p$-local finite spectrum $F$ such that $F\otimes X$ is $\pi$-finite (bounded, and the homotopy groups are finite groups). We say $X$ has fp-type $n$, if $n=\min\{k-1\mid F\otimes X~\text{is}~\pi\text{-finite},F~\text{is fp-type}~k\}$. A spectrum $X$ is said to be pure fp-type $n$, if moreover $F_{*}X$ is a free finitely generated $P(v)$-module for some finite CW-spectrum of type $n$, with $v_{n}$-map $v:\Sigma^{d}F\to F$. In his 2000 Oberwolfach lecture [ROG00], Rognes introduced the term ”redshift”, and the redshift problem is stated as follows.
@@
In chromatic homotopy theory, $v_{n}$ is detected by Morava K-theory, $K(n)$. One could use Morava K-theory to define the height for spectra, although, in general, this is not a very nice concept. We say an $\mathbb{E}_{\infty}$-ring $R$ is of height $n$, if $L_{K(n)}R\not\simeq 0$, and $L_{K(n+1)}R\simeq 0$. Hahn [HAH22] showed that for an $\mathbb{E}_{\infty}$-ring, if $L_{K(i)}R\simeq 0$, then $L_{K(j)}R\simeq 0$, for all $j\geq i$. Thus, in this case, the height is well defined. Let $R$ be a ring spectrum. So, one might define the height of $R$ to be the number $n$ such that $K(n)_{*}R\neq 0$ and $K(m)_{*}R=0$ for $m>n$.
@@
The conjecture is true for $R$ any $\mathbb{E}_{\infty}$-ring, by [YUA21], [BSY22], [CMN+24], [LMM+24]. For the relation of the first conjecture to $T(n)$-localized theory, see Theorem 4.5. The second conjecture is more related to $K(n)$-localized theory. As known, for $n\geq 2$, [BHL+23] shows the telescope conjecture fails, using many ingredients in algebraic K-theory and things related to these two conjectures.
@@
Let $R$ be an $\mathbb{E}_{1}$-ring spectrum of (pure) fp type $n$. Does $\operatorname{TR}(R,p)$ have fp type $n+1$.
The method to prove Conjecture 1.4, is first to construct a nice algebra map $R\to S$, yielding an algebra map $L_{K(n+1)}\operatorname{K}(R)\to L_{K(n+1)}\operatorname{K}(S)$. If one can show that $L_{K(n+1)}\operatorname{K}(S)\not\simeq 0$, then one could conclude that the source is not 0.
The K-theory of truncated polynomial algebras over perfect fields were studied in [HM97a]. Note that $\mathbb{F}_{p}[x]/x^{e}$ is just $\operatorname{BP}\langle-1\rangle[x]/x^{e}$. Thus it is natural to consider the K-theory of $\operatorname{BP}\langle n\rangle[x]/x^{e}$, which are chromatic higher level cases of $\mathbb{F}_{p}[x]/x^{e}$. The goal of this paper is to study Problem 1.2, Conjecture 1.3, 1.4, Question 1.5, 1.6, for $R=\operatorname{BP}\langle n\rangle[x]/x^{e}$, and other rings related to that. In particular, in the proof of Proposition 4.9 and [HM97a, Theorem A], shows that $\operatorname{K}(\mathbb{F}_{p}[x]/x^{e})$ is not an fp spectrum of type 0. And Proposition 6.4, shows that $\operatorname{K}((\ell[x]/x^{e})^{\wedge}_{p})^{\wedge}_{p}$ is not an fp spectrum of type $2$.
@@ 1.2. Main results
The chromatic redshift problem 1.2 does not assume regularity hypothesis, i.e. does not require $R$ to be ”regular” in some sense. In this paper, we show that a regularity hypothesis is in fact necessary, see Theorem 1.14. The truncated Brown–Peterson spectra may be regarded, in a suitable sense, as regular ring spectra, in light of Theorem 1.12 and 1.13.
Remark 1.9.
The term ”higher local fields” first appeared in [ROG05] (the chromatic height one case). The author suspected that the truncated Brown–Peterson spectra should be thought as the valuation rings of ”higher local fields”, a phrase that also appears in the title of this paper. In other words, the truncated Brown–Peterson spectra may be thought as ”higher local number rings”. After the completion of this work, Rognes and his collaborators defined the notion ”higher local number ring”, see [ROG25]. This definition appears to be closely related to Theorem 1.14.
@@
See Convention 2.12, 2.19. Any of them is a finite $\operatorname{BP}\langle n\rangle$-module for some $n$, thus we can view them as finite algebras over $\operatorname{BP}\langle n\rangle$ (the ring of integers of a height $n$ local field). Moreover $(\operatorname{BP}\langle n\rangle[x]/x^{e})^{\wedge}_{p},(\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\operatorname{MU}[x]/x^{e})^{\wedge}_{p}$ and $(\Lambda_{\operatorname{BP}\langle n\rangle}(x))^{\wedge}_{p},(\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}\Lambda_{\operatorname{MU}}(x))^{\wedge}_{p}$ are fp spectra of type $n$, and the spectrum $(\operatorname{BP}\langle n\rangle/v_{n}^{e})^{\wedge}_{p}$ is an fp spectrum of type n1n-1.
@@
induces isomorphism on large degrees. In particular, if $F$ is of type $n+2$, the above also holds. If one replaces $\operatorname{BP}\langle n\rangle$ by $\operatorname{BP}\langle n\rangle\otimes\mathbb{S}_{W(k)}$ the statement holds still. Here, $k$ is a perfect field of characteristic $p$, $\mathbb{S}_{W(k)}$ is the spherical Witt vectors.
@@
$\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])\to L_{2}^{f}(\operatorname{TC}(\operatorname{BP}\langle 1\rangle[x])),$
$\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}\to L_{2}^{f}(\operatorname{K}(\operatorname{BP}\langle 1\rangle[x])^{\wedge}_{p}).$
In the classical case, algebraic K-theory has the $\mathbb{A}^{1}$-invariance property for regular rings. This result shows that algebraic K-theory has $\mathbb{A}^{1}$-invariance property for regular ring spectra in the following sense. If $R$ is a regular ring spectrum such that $F\otimes\operatorname{K}(R)$ is bounded for a certain ($p$-local) finite spectrum $F$, then $F\otimes\operatorname{K}(R[x])$ is still bounded.
By decomposing $\operatorname{THH}(\mathbb{S}[x]/x^{e})$ into a sum of $S^{1}$-spectra, and analyzing the Frobenius map, we found the following theorem. Actually, the base case is Proposition 4.9.
@@
Let $F$ be any $p$-local type $n+2$ finite complex. Let $W$ denote the fiber of Frobenius map of $F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)$. The following statements are true.
  1. (1)
    If $W\not\simeq 0$, then the fiber of the Frobenius
@@
    is not bounded above, where $e>1$. In other words, $F$-Segal conjecture for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is true if and only if $W\simeq 0$.
  2. (2)
    If $F$ is a sum of generalized Moore complexes $\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}},\ldots v_{n+1}^{i_{n+1}})$, then $W\not\simeq 0$.
  3. (3)
    In fact, for $F$ a $p$-local type $n+2$ finite complex, $W\not\simeq 0$.
  4. (4)
    If $e>1$, the spectrum $F\otimes\operatorname{TR}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ is not bounded. In other-words, the Lichtenbaum–Quillen property defined in Definition 1.7 fails for $\operatorname{BP}\langle n\rangle[x]/x^{e}$.
@@
$\operatorname{THH}(\operatorname{MU}[x]/x^{e})\to\operatorname{THH}(\operatorname{MU}[x]/x^{e})^{tC_{p}},$
$\operatorname{THH}(\operatorname{BP}[x]/x^{e})\to\operatorname{THH}(\operatorname{BP}[x]/x^{e})^{tC_{p}},$
@@ 1.3. Future work
It seems that one could use the method presented in this paper to investigate some other examples of the Segal conjecture, by using the cyclic decomposition, i.e. decomposition of $\operatorname{THH}(\mathbb{S}[M])$ or $\operatorname{THH}(\mathbb{S}[\Omega Y])$, to carry out more computations, where $M$ is a commutative monoid, $Y$ is a space. For instance, in a forthcoming work, we aim to compute $\operatorname{TC}(\operatorname{BP}\langle n\rangle[x,y]/(xy))$ at low heights. Similarly, on could replace $\operatorname{BP}\langle n\rangle$ by $K(n)$ or $k(n)=\tau_{\geq 0}K(n)$.
@@
Question 1.16 ([HLS24, Question 1.23]).
Does mod $(p,v_{1},\ldots,v_{n+1})$ syntomic cohomology for height $n$ ring spectra, factor though its mod $(p^{i_{0}},v_{i}^{i_{1}},\ldots,v_{n}^{i_{n}})$ reduction for some sequence $i_{0},\ldots,i_{n}$ ?
@@ 1.4. Structure of the paper
In Section 2, we use several $\mathbb{E}_{\infty}$-ring maps to construct several $\mathbb{E}_{3}$-algebras over some base ring spectra other than $\mathbb{S}$, and we also use the tower of $\mathbb{E}_{\infty}$-algebras in [HW22, §2] to get an $\mathbb{E}_{3}$-algebra $\operatorname{BP}\langle n\rangle/v_{n}^{e}$. All these algebras are finite algebras over $\operatorname{BP}\langle n\rangle$, see Proposition 2.28. In Section 3, we use several $\mathbb{E}_{3}$-algebra maps to prove Conjecture 1.4, and to prove Theorem 1.11. In Section 4, we use the descent functor $\operatorname{desc}_{\mathbb{F}_{p}}$ to prove the Segal conjecture, and we use the decompositions of $\operatorname{THH}(\mathbb{S}[x]/x^{e}),\operatorname{THH}(\mathbb{S}[\Omega S^{3}])$ to prove the failure of the Segal conjecture for some ring spectra. The primary goal is to prove Theorem 1.12, 1.13, 1.14. In Section 5, we use the evenly free map $\operatorname{THH}(\operatorname{MU}[x])\to\operatorname{MU}[x]$ to compute the $E_{2}$-page of the descent spectral sequence which computes $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e};\mathbb{F}_{p}[x]/x^{e})$. We show that the spectral sequence degenerates at the $E_{2}$-page in some cases. See Proposition 5.4. In Section 6, we use the decomposition of $\operatorname{THH}(\mathbb{S}[x]/x^{e})$ try to carry out some computations for $V(2)_{*}\operatorname{TC}(\ell[x]/x^{e})$. We prove Proposition 6.4.